OpenStudy (princeevee):

solve for x 3/4x=9 1/4 4/3 4 -3/4

OpenStudy (firekat97):

your goal is to isolate x, so firstly you'd want to bring it into the numerator position, any ideas on how you may be able to do that?

OpenStudy (princeevee):

no ;-;

OpenStudy (anonymous):

is the problem \[\frac{ 3 }{4 }x = 9 \] or is it \[\frac{ 3 }{ 4x } = 9\]?

OpenStudy (firekat97):

oh okay so what we can do firstly is multiply both sides by 4x... this way, we're not changing the equation and we can get x in the numerator position on the RHS (as x will cancel on the LHS so like.. \[\frac{ 3 }{ 4x } . 4x = 9 . 4x\]

OpenStudy (firekat97):

so try simplifying that and post up what you get

OpenStudy (princeevee):

the first one lily

OpenStudy (firekat97):

@princeevee does that make sense?

OpenStudy (firekat97):

oh wait lol I got the question wrong nvm

OpenStudy (anonymous):

oh then \[\frac{ 3 }{ 4 }x*\frac{ 4 }{ 3 } = 9* \frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

@FireKat97 :)

OpenStudy (princeevee):

so B?

OpenStudy (anonymous):

x= 12 , but ya B would match your options.

OpenStudy (princeevee):

k